# Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge

### v1

@@ 8 Limitations
While this study provides valuable insights into cross-cultural knowledge probing in LLMs, it is essential to acknowledge several limitations. Firstly, the food domain knowledge dataset utilized in this research is sourced from Wikidata, which may not offer comprehensive coverage. For example, the dish soy sauce chicken may only include the ingredient chicken meat while lacking the inclusion of soy sauce. Moreover, ingredient descriptions are not always detailed. For instance, the Wikidata gold label might be oil when the recipe requires a specific type of oil, such as sesame oil. This inconsistency underscores the motivation behind our mWS metric. Furthermore, aside from well-known dishes, certain recipes lack standardization and may vary depending on individual preferences and cooking styles, posing challenges to precise probing. Additionally, the fuzzy-match metric mWS, introduced in this study, relies on Fasttext for obtaining object representation vectors. However, for certain objects in Chinese and Korean, zero vectors may result, rendering similarity calculation impossible. Lastly, we employ manually crafted templates in this paper. However, research has shown that sampling templates from large corpora can also enhance knowledge-probing evaluation. This aspect is deferred to future work. Despite our endeavors to construct comprehensive multilingual and multicultural knowledge repositories, the availability of aligned cross-cultural knowledge remains limited in multilingual settings. This constraint presents challenges in exploring the interaction between language and culture.
Appendix A FmLAMA examples
@@
Particularly, in each instance, $la=\mathrm{LANG}\left(na\right)=\mathrm{LANG}\left(pa\right)=\mathrm{LANG}\left(ma\right)$. For a dish with the same $url$, there may be several instances with different languages. There are totally of 33,601 entries in FmLAMA.
@@ Appendix B Probing Templates
Appendix C Probing results in each FmLAMA-la
Figure 5: Probing results with code-switching settings in each cultural group, w/o and w/ signify the absence or presence of cultural background information in the prompt, respectively.
Figure 6: Examples of FmLAMA.
@@ Appendix C Probing results in each FmLAMA-la
Besides the multilingual LLMs discussed in §5.1, we also probe XLM-RoBERTa Conneau et al. (2020) here. [^10] Indeed, we probe XLM-RoBERTa on FmLAMA-en as well, yielding very poor probing results. Nonetheless, although XLM-RoBERTa’s performance remains subpar in other languages, it shows some relatively positive outcomes. Furthermore, in addition to Hebrew, we configure probing for monolingual LLMs in Arabic, Russian, Korean, and Chinese, including asafaya/bert-base-arabic Safaya et al. (2020), DeepPavlov/rubert-base-cased Kuratov and Arkhipov (2019), kykim/bert-kor-base, klue/bert-base Park et al. (2021) and bert-base-chinese.
@@
The probing results with prompts in the other five languages (Arabic, Hebrew, Russian, Korean, and Chinese) on the corresponding filtered sub-datasets (FmLAMA-ar, FmLAMA-he, FmLAMA-ru, FmLAMA-ko, and FmLAMA-zh) are depicted in Table 6 through Table 10, respectively. Because certain objects in Chinese and Korean have representation vectors that result in all zeros when obtained through Fasttext, calculating cosine similarity was not feasible. Consequently, mWS evaluation was not conducted for prompts in Chinese and Korean prompts. Overall, irrespective of the language used for probing, LLMs still exhibit a relatively strong familiarity with knowledge in the food domain within American culture. However, they may show a slight preference for certain cultural knowledge; for example, when probing in Arabic, LLMs may show a better capability in probing knowledge related to Iranian groups. Additionally, the probing capability of monolingual LLMs may not necessarily surpass that of multilingual LLMs.
Appendix D Cultural results for Code-switching probing
For a more detailed comparison of the knowledge probing abilities of LLMs across various cultural groups, we evaluate the probing results for each cultural group individually. Figure 5 shows the average probing results for the three multilingual LLMs in each cultural group, from which we find: (1) English-dominant prompt settings generally outperform those where Chinese is the primary language. (2) En_en prompt setting excels in probing food knowledge for U.K., U.S., India, and China cultures, while proficiency in the ZH_zh prompt setting is also notable for the above cultural groups except India. (3) In code-switching settings, the detection abilities of EN_zh relative to EN_en and ZH_en relative to ZH_zh both notably decrease, with the latter showing a more pronounced decline. (4) Whether in code-switching mode or not, introducing relevant cultural backgrounds in the prompts aids in the detection of cross-cultural knowledge, manifested in overall probing results and the probing ability within almost every cultural group. The more detailed probing results for each multilingual LLM across various cultural groups can be found in Tables 11 and 12.
Appendix E Data distribution of 175 dishes
The figure 8 illustrates the data distribution of the filtered sub-dataset used in the language analysis in §5.2. It encompasses both continent-level and country-level data distributions.
(a) Continent-level distribution
(b) Country-level distribution
Figure 8: Data distribution for the filtered sub-dataset comprising 175 dishes.
Appendix F Decoder-only LLM probing details
In this section, We present the experimental setup for knowledge probing with decoder-only LLMs, and showcase their fine-grained probing results across various cultural groups.
F.1 Experiential setup
Decoder-only LLMs
: LlaMa2-chat Touvron et al. (2023), a collection of pre-trained and fine-tuned chat models; Vicuna Chiang et al. (2023), a chat assistant trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT; and gpt-3.5-turbo* OpenAI (2023), a GPT-3.5 model fine-tuned on human instructions using Reinforcement Learning with Human Feedback (RLHF).
Prompt construction
Decoder-only LLMs, in contrast to encoder-only LLMs and encoder-decoder LLMs, focus solely on generating text based on contextual information provided to them. Considering these factors, the probing task involving predicting the probability of the [MASK] token during decoder-only LLM probing will no longer be applicable. We extend the initial template with fill-in instructions, such as “The food [X] has the ingredients of []. Please fill in the sentence.” We only select the best-performing template for extension and probing of the decoder-only LLMs here. Taking into account the diversity of responses from chat LLMs and the ease of evaluating the probing results, we also consider employing an alternative fill-in instruction: “Please complete the sentence with only one entity”. The final results only showcase the probing outcomes under the best-performing settings.
Evaluation
For each dish instance, we parse the text generated by the decoder-only LLMs to obtain the final predicted object list. This object list is not derived from a candidate object set, resulting in greater diversity. Therefore, before conducting the matching evaluation, we lemmatize each object in both the golden object list and the predicted object list to obtain the lemmatized format of each word, such as potatoes $\rightarrow$ potato. We utilize accuracy (ACC) as the metric here, employing a forgiving form of absolute matching to assess result accuracy. Specifically, for each dish instance, if any predicted object matches any object in the golden object list, we consider the prediction for that dish to be correct.
F.2 Cultural probing results
Table 13 show the probing results in each cultural group, including monolingual prompts and code-switching prompts. In monolingual English prompts, all decoder-only LLMs exhibit superior probing performance for knowledge related to the cultural groups of Italy, the U.S., and France. When using English-Chinese code-switching prompts, the overall probing performance tends to decrease, but the best-performing cultural groups remain largely unchanged. However, there are some improvements in the probing rankings for the China group in Llama2-7b-chat-hf and the U.K. group in vicuna-13b. In the case of monolingual Chinese prompts, the cultural groups with the best probing performance are primarily concentrated in China, the U.K., and the U.S. Similarly, when using Chinese-English code-switching prompts, the overall probing performance tends to decrease.
Origin Count Bb-ar Bl-ar mB-c mB-u XRb XRl mT5 Avg.
Italy 76 (13.3%) 2.16±0.95 2.55±0.97 3.52±1.73 2.55±1.25 1.33±0.63 1.30±0.43 3.35±0.71 2.39
U.S. 54 (9.5%) 3.07±1.45 6.01±3.35 4.74±2.46 4.32±4.47 2.88±2.46 2.26±1.05 6.17±0.97 4.21
Turkey 52 (9.1%) 4.59±2.30 5.97±4.08 4.58±1.37 4.41±3.98 2.50±1.14 1.93±0.11 3.18±0.90 3.88
Japan 44 (7.7%) 3.02±1.11 2.86±1.01 3.76±1.54 2.80±0.56 1.69±0.93 1.30±0.10 2.18±0.77 2.52
France 37 (6.5%) 5.25±2.06 5.87±4.09 4.92±3.30 3.98±3.64 4.28±4.26 3.67±2.76 6.17±2.40 4.88
U.K. 26 (4.6%) 6.78±2.88 8.97±6.33 8.53±5.48 6.02±6.67 2.77±3.28 2.18±1.86 8.84±5.02 6.30
Mexico 20 (3.5%) 2.95±0.40 5.44±2.22 4.52±2.85 5.74±1.82 2.08±0.34 1.88±0.27 3.24±1.27 3.69
India 18 (3.2%) 3.67±3.17 6.41±3.58 9.02±7.34 4.44±5.49 1.90±0.27 1.79±0.13 5.55±2.93 4.68
Germany 13 (2.3%) 6.92±3.83 9.67±5.54 8.03±4.78 6.05±4.50 1.45±0.72 1.49±0.65 5.41±3.72 5.57
China 13 (2.3%) 3.25±2.55 2.91±2.51 4.44±3.22 4.33±4.76 1.93±0.16 1.76±0.10 4.04±2.01 3.24
Iran 11 (1.9%) 7.74±5.56 11.09±5.02 6.61±4.79 4.61±5.41 4.04±4.51 4.10±4.54 6.02±4.34 6.32
Greece 11 (1.9%) 3.31±0.85 4.98±4.35 4.16±3.98 4.66±4.18 1.79±1.15 1.72±0.89 4.04±1.80 3.52
Spain 10 (1.8%) 6.03±2.68 4.75±3.35 6.16±4.09 3.04±2.87 3.46±1.32 2.79±0.27 5.53±1.39 4.54
Russia 10 (1.8%) 3.36±0.63 2.32±0.49 1.93±0.27 1.82±0.32 2.05±0.41 1.91±0.21 2.47±0.48 2.27
Others 176 (30.8%) 4.03±0.97 3.54±1.57 4.30±1.40 2.83±1.72 2.74±1.19 2.31±0.15 4.56±1.59 3.47
ALL 571 (100.0%) 3.95±1.31 4.66±2.36 4.73±2.03 3.61±2.66 2.48±1.46 2.10±0.59 4.53±1.37 3.72
@@ F.2 Cultural probing results
Bb-ar Bl-ar mB-c mB-u XRb XRl mT5 Average
Italy 0.3147±0.02 0.2786±0.05 0.2917±0.04 0.2383±0.02 0.1991±0.02 0.1930±0.00 0.3120±0.03 0.2611
U.S. 0.2943±0.03 0.3057±0.05 0.3087±0.02 0.2648±0.05 0.2208±0.05 0.2072±0.02 0.3264±0.02 0.2754
Turkey 0.2996±0.03 0.2754±0.07 0.2787±0.04 0.2432±0.03 0.1941±0.03 0.1866±0.01 0.2867±0.02 0.2520
Japan 0.2977±0.02 0.2416±0.04 0.2685±0.02 0.2409±0.03 0.1795±0.02 0.1729±0.01 0.2791±0.03 0.2400
France 0.3360±0.02 0.3221±0.05 0.3032±0.04 0.2877±0.03 0.2255±0.08 0.2103±0.05 0.3333±0.04 0.2883
U.K. 0.3125±0.05 0.3418±0.06 0.3632±0.06 0.2736±0.09 0.2270±0.07 0.2103±0.03 0.3681±0.04 0.2995
Mexico 0.3273±0.02 0.2866±0.04 0.3036±0.03 0.2923±0.04 0.2039±0.04 0.2043±0.04 0.3036±0.04 0.2745
India 0.2707±0.03 0.2726±0.07 0.3030±0.04 0.2261±0.06 0.1950±0.02 0.1814±0.01 0.2808±0.03 0.2471
Germany 0.3106±0.03 0.3321±0.06 0.3275±0.05 0.2758±0.05 0.2048±0.03 0.1923±0.00 0.3113±0.02 0.2792
China 0.3006±0.04 0.2658±0.04 0.3366±0.03 0.2854±0.06 0.2035±0.02 0.2000±0.01 0.3037±0.05 0.2708
Iran 0.3714±0.05 0.3420±0.06 0.3370±0.05 0.2687±0.05 0.2291±0.07 0.2286±0.07 0.3289±0.04 0.3008
Greece 0.3068±0.01 0.3067±0.06 0.3260±0.05 0.2693±0.03 0.2270±0.01 0.2306±0.02 0.3147±0.04 0.2830
Spain 0.3930±0.03 0.3049±0.06 0.3487±0.06 0.3124±0.03 0.2377±0.03 0.2395±0.03 0.3418±0.05 0.3111
Russia 0.3353±0.01 0.3068±0.04 0.2885±0.03 0.2784±0.01 0.2264±0.05 0.2183±0.03 0.3157±0.03 0.2813
Others 0.2989±0.01 0.2720±0.05 0.2932±0.04 0.2452±0.01 0.2049±0.04 0.1939±0.01 0.3001±0.02 0.2583
ALL 0.3078±0.02 0.2854±0.05 0.2999±0.03 0.2549±0.02 0.2065±0.04 0.1972±0.02 0.3083±0.02 0.2657

### v3

@@ 9 Limitations
While this study provides valuable insights into cross-cultural knowledge probing in LLMs, it is essential to acknowledge several limitations. Firstly, the food domain knowledge dataset utilized in this research is sourced from Wikidata, which may not offer comprehensive coverage. For example, the dish soy sauce chicken may only include the ingredient chicken meat while lacking the inclusion of soy sauce. Moreover, ingredient descriptions are not always detailed. For instance, the Wikidata gold label might be oil when the recipe requires a specific type of oil, such as sesame oil. This inconsistency underscores the motivation behind our mWS metric. Furthermore, aside from well-known dishes, certain recipes lack standardization and may vary depending on individual preferences and cooking styles, posing challenges to precise probing. Additionally, the fuzzy-match metric mWS, introduced in this study, relies on Fasttext for obtaining object representation vectors. However, for certain objects in Chinese and Korean, zero vectors may result, rendering similarity calculation impossible. Furthermore, we employ manually crafted templates in this paper. However, research has shown that sampling templates from large corpora can also enhance knowledge-probing evaluation. This aspect is deferred to future work. Despite our endeavors to construct comprehensive multilingual and multicultural knowledge repositories, the availability of aligned cross-cultural knowledge remains limited in multilingual settings. This constraint presents challenges in exploring the interaction between language and culture.
ACKNOWLEDGMENTS
This research is supported by the project of Shenzhen Science and Technology Research Fund (Fundamental Research Key Project Grant No. JCYJ20220818103001002), Shenzhen Science and Technology Program (Grant No. ZDSYS20230626091302006), Key Project of Shenzhen Higher Education Stability Support Program (Grant No. 2024SC0009).
Appendix A FmLAMA details
Datasets Format Topic Construction method Size
FmLAMA (ours) Triplet knowledge Food domain Automatically constructed 33600
A.1 Examples
@@
Particularly, in each instance, $la=\mathrm{LANG}\left(na\right)=\mathrm{LANG}\left(pa\right)=\mathrm{LANG}\left(ma\right)$, which indicates the descriptive language for all attributes remains consistent. For a dish with the same $url$, there may be several instances with different languages. Table 5 contrasts our dataset with prior cultural knowledge collections.
A.2 Statistics
There are a total of 33,600 dishes in FmLAMA. We count the number of dishes corresponding to each language and country in the dataset, as shown in Figures 6 and 7. Figure 6(a) shows the top 20 languages ranked by the number of dishes, with English (en) having the highest count at 2804, followed by Spanish (es), French (fr), and Japanese (ja). Figure 6(b) displays the distribution of dish counts across different intervals for all languages, indicating that most languages have fewer than 200 dishes. Similarly, Figure 7(a) shows the top 20 countries ranked by the number of dishes, with Italy having the highest count at 2975, followed by France, America, and Japan. Figure 7(b) shows the distribution of dish counts across various intervals for all countries, indicating that only 8 countries have more than 200 dishes. Figure 8 shows the dish statistics by ingredient count. The majority of dishes contain only one ingredient, totaling 18,546. Dishes with two ingredients are the second most common, with a count of 6,426. The counts decrease as the number of ingredients increases, with dishes containing eleven ingredients being the least common, at 32.
(a) The Top20 languages ranked by the dishes count
(b) Languages distributed across various dish count ranges
Figure 6: Dish statistics by languages
(a) The Top20 countries ranked by the dishes count
(b) Countries distributed across various dish count ranges
Figure 7: Dish statistics by countries
Figure 8: Dish statistics by ingredient count
@@ B.1 Prompt illustration
B.2 Language comparison
B.2.1 Geographic differences
In this section, we discuss the representative countries where each selected language is spoken and their geographic distribution: (1) English: Mainly spoken in U.S., Canada, Australia, Guyana, etc., spanning North and South America, Europe, Australia, and Africa. (2) Chinese: Mainly spoken in China, Singapore, etc., covering huge area of Asia. (3) Arabic: Mainly spoken in Saudi Arabia, Egypt, United Arab Emirates, etc., covering the Arabian Peninsula, North Africa, and some sub-Saharan African countries. (4) Korean: Mainly spoken in South Korea, North Korea. (5) Russian: Mainly spoken in Russia, Belarus, Kazakhstan, etc., covering Eastern Europe, Central Asia, and Northern Asia. (6) Hebrew: Mainly spoken in Israel, with Hebrew-speaking communities worldwide. These languages represent cultural diversity, being spoken on different continents by groups with rich and distinct cultural backgrounds, demonstrating the broad geographic and cultural diversity represented by these languages.
B.2.2 Grammatical differences
In addition to geographical differences in usage, these six languages also display the following grammatical and reading distinctions: (1) Word Order: English and Chinese follow an Subject-Verb-Object word order, Arabic typically uses Verb-Subject-Object, Korean follows Subject-Object-Verb, while Russian and Hebrew are relatively flexible but generally adhere to Subject-Verb-Object. (2) Tense: English, Arabic, Russian, and Hebrew possess complex tense systems. Chinese lacks a strict tense structure, whereas Korean expresses tense through verb conjugation. (3) Gender: Arabic, Russian, and Hebrew have grammatical gender, while English distinguishes gender only in pronouns. Chinese and Korean do not have grammatical gender. (4) Articles: English, Arabic, and Hebrew have an article system, whereas Chinese, Korean, and Russian do not use articles. (5) Reading Direction: English, modern Chinese, modern Korean, and Russian are read from left to right, while Arabic and Hebrew are read from right to left. Traditional Chinese and traditional Korean are read vertically from top to bottom, with columns arranged from right to left. (6) Case System: Arabic and Russian have complex case systems, while Korean employs particles to indicate grammatical relations. English, Chinese, and Hebrew do not use case inflections.
Appendix C Manual Evaluation Setup
C.1 GPT-4o simulator
The evaluation prompts for the GPT-4o simulator are shown in Table 6.
C.2 Human evaluator
Figure 10 shows the human evaluation website. The Inter-Rater Reliability (IRR) and the average evaluation time for the three evaluators from each country group are provided in Table 7.
Instructions
You are a professional evaluator. I will provide you with a dish, its reference ingredient label sourced from Wikidata, and a list of predicted ingredients. Your task is to evaluate whether each predicted ingredient is appropriate for the respective dish.
### Evaluation Criteria:
For each predicted ingredient, mark it as Correct, Maybe, or Incorrect using the following guidelines:
Correct: The ingredient meets one of these: Direct Match: It is explicitly listed in the reference ingredient label from Wikidata. Substitutability: It can replace a specific ingredient in the reference label during cooking. Missing Traditional Ingredient: It is traditionally or commonly used in the dish, but not listed in the reference label. Maybe: The ingredient could be used in some variations of the dish, but its use is uncommon or ambiguous. Incorrect: The ingredient is rarely or never used in the dish. ### Output Format:
Please directly return the evaluation as a Python dictionary with JSON format where the key is the predicted ingredient, and the value is a tuple consisting of:
1. The evaluation ("Correct", "Maybe", "Incorrect")
2. The reason ("Direct Match", "Substitutability", "Missing Traditional Ingredient", or a brief explanation)
### Example:
Input:
Dish: Spaghetti Bolognese
Reference ingredient label: spaghetti, ground beef, onion, garlic, Carrot
Predicted ingredients: spaghetti, basil, chicken, tomato, beef
Output
    ’ground beef’: (’Correct’, ’Substitutability’),
    ’meat’: (’Maybe’, ’Ambiguous’),
    ’beef’: (’Correct’, ’Substitutability’),
    ’pasta’: (’Incorrect’, ’Pasta is not part of the sauce itself’),
    ’turkey meat’: (’Maybe’, ’It can be used in some variations’),
Table 6: GPT-4o simulator: Evaluation Instructions
Figure 10: Human evaluation website
Group Dishes IRR Time(m)
Italy 215 0.60 74.93
U.S. 285 0.48 104.25
Turkey 98 0.78 32.05
Japan 186 0.64 49.25
France 175 0.63 71.83
U.K. 83 0.46 30.33
Mexico 57 0.44 22.25
India 132 0.58 40.25
Germany 57 0.69 20.03
China 97 0.59 48.50
Iran 21 0.61 13,33
Greece 21 0.35 11.30
Spain 95 0.57 41.50
Russia 27 0.51 26.50
Mean 111 0.57 -
Table 7: Human evaluator: IRR and average evaluation time for the three evaluators from each country group.
Appendix D Supplementary Experiments and Analyses
D.1 Probing results in each FmLAMA-la
@@
In addition to the multilingual LLMs discussed in §5, configure probing for monolingual, encoder-only LLMs in Arabic, Russian, Korean, and Chinese, including asafaya/bert-base-arabic Safaya et al. (2020), DeepPavlov/rubert-base-cased Kuratov and Arkhipov (2019), kykim/bert-kor-base, klue/bert-base Park et al. (2021) and bert-base-chinese.
@@
The probing results with prompts in the other five languages (Arabic, Chinese, Hebrew, Korean, and Russian) on the corresponding filtered sub-datasets (FmLAMA-ar, FmLAMA-zh, FmLAMA-he, FmLAMA-ko, and FmLAMA-ru) are depicted in Table 9101112, and 13, respectively. Because certain objects in Chinese and Korean have representation vectors that result in all zeros when obtained through Fasttext, calculating cosine similarity was not feasible. Consequently, mWS evaluation was not conducted for prompts in Chinese and Korean prompts.
Origin Encoder-only LLMs Encoder-Decoder LLMs Decoder-only LLMs Avg.
Bb Bl mB mT5 T5 Qwen2 Llama2 Llama3
Italy 0.7393±0.01 0.7425±0.01 0.7438±0.00 0.6865±0.02 0.6579±0.01 0.7749±0.01 0.7658±0.01 0.7869±0.01 0.7372
U.S. 0.7748±0.01 0.7763±0.01 0.7815±0.01 0.7071±0.02 0.7178±0.03 0.7968±0.01 0.7839±0.02 0.8107±0.01 0.7686
Turkey 0.7357±0.01 0.7461±0.01 0.7545±0.01 0.7015±0.02 0.6703±0.01 0.7681±0.01 0.7571±0.01 0.7830±0.01 0.7395
Japan 0.7113±0.00 0.7144±0.00 0.7218±0.01 0.6788±0.02 0.6442±0.01 0.7446±0.00 0.7446±0.01 0.7516±0.00 0.7139
France 0.7398±0.01 0.7520±0.02 0.7513±0.01 0.6755±0.02 0.6642±0.01 0.7941±0.01 0.7796±0.01 0.7846±0.00 0.7426
U.K. 0.7460±0.01 0.7500±0.01 0.7550±0.00 0.6697±0.02 0.6745±0.02 0.7678±0.01 0.7701±0.01 0.7790±0.01 0.7390
Mexico 0.7273±0.01 0.7313±0.01 0.7291±0.01 0.7171±0.01 0.6734±0.01 0.7605±0.01 0.7752±0.02 0.7852±0.02 0.7374
India 0.7477±0.01 0.7523±0.00 0.7568±0.00 0.6998±0.01 0.6619±0.01 0.7867±0.00 0.7944±0.01 0.7935±0.01 0.7491
Germany 0.7348±0.01 0.7305±0.03 0.7493±0.00 0.6755±0.01 0.6535±0.01 0.7835±0.01 0.7769±0.02 0.7849±0.01 0.7361
China 0.7273±0.01 0.7319±0.01 0.7317±0.01 0.7024±0.01 0.7054±0.01 0.7698±0.01 0.7593±0.01 0.7793±0.01 0.7384
Iran 0.7064±0.01 0.7013±0.01 0.7310±0.01 0.7042±0.02 0.6686±0.01 0.7905±0.03 0.7914±0.04 0.8026±0.01 0.7370
Greece 0.7488±0.01 0.7618±0.02 0.7566±0.02 0.7190±0.03 0.6369±0.00 0.8137±0.04 0.7739±0.01 0.8247±0.01 0.7544
Spain 0.7402±0.00 0.7465±0.01 0.7472±0.00 0.6883±0.01 0.6391±0.01 0.7748±0.01 0.7663±0.01 0.7765±0.01 0.7349
Russia 0.7467±0.01 0.7571±0.01 0.7573±0.00 0.7216±0.02 0.6710±0.01 0.7716±0.01 0.7680±0.01 0.7721±0.00 0.7457
ALL 0.7419±0.00 0.7466±0.01 0.7507±0.00 0.6924±0.01 0.6723±0.01 0.7778±0.01 0.7711±0.01 0.7859±0.00 0.7423
Corr. 0.35 0.59 0.38 0.83 0.57 0.86 0.71 0.96 -
Table 8: Probing performance comparison with English prompts and FmLAMA-ar sub-dataset. The results are evaluated using mWS with BERT base. Corr. denotes the correlation of mWS evaluations using BERT Base and FastText for each LLM, which reveal a strong correlation on decoder-only LLMs,
Origin Count Bb-ar mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 76 (19.2%) 2.56±0.69 3.07±1.02 5.12±1.20 5.13±3.42 3.30±1.13 4.05±0.81 4.03
U.S. 54 (13.7%) 5.29±1.18 7.22±3.91 5.68±1.53 9.41±4.12 3.93±1.00 5.95±2.38 5.09
Turkey 52 (13.2%) 4.07±0.88 6.66±3.09 3.30±0.53 9.20±5.71 3.33±1.04 9.54±3.48 5.47
Japan 44 (11.11%) 2.69±1.50 2.56±0.50 2.13±0.78 9.91±3.73 3.75±1.12 5.43±1.62 5.20
France 37 (9.4%) 7.51±2.43 6.50±4.55 4.51±1.21 7.16±3.23 5.71±0.82 4.63±1.77 4.70
U.K. 26 (6.6%) 5.80±2.15 9.17±6.85 6.72±1.67 11.77±6.35 5.72±4.86 8.62±6.52 5.75
Mexico 20 (4.6%) 3.52±1.57 7.37±1.01 3.43±0.61 3.60±0.55 2.38±0.48 3.72±1.68 2.82
India 18 (3.2%) 4.39±1.09 8.86±5.19 3.08±1.17 10.82±9.44 3.04±1.56 3.03±1.38 7.29
Germany 13 (3.3%) 7.39±2.52 10.65±4.95 5.87±2.04 9.53±4.72 2.75±2.12 7.67±3.29 4.69
China 13 (3.3%) 2.07±0.53 4.86±4.36 2.43±0.93 6.79±2.17 2.75±0.70 6.54±2.51 4.58
Iran 11 (2.8%) 9.30±4.85 8.19±5.41 7.38±2.56 11.23±7.82 6.31±2.85 6.67±5.02 5.90
Greece 11 (2.8%) 3.12±0.98 5.21±3.26 4.25±1.23 9.25±4.78 3.13±1.09 4.54±3.95 4.63
Spain 10 (2.5%) 5.84±1.24 4.13±2.32 8.01±1.35 12.74±3.84 3.41±1.18 6.90±4.08 7.34
Russia 10 (2.5%) 3.62±0.80 3.13±0.79 2.48±0.26 4.20±0.98 4.88±2.78 4.76±0.94 4.34
ALL 395 (100.0%) 4.41±0.35 5.80±2.56 4.48±0.60 8.24±3.65 3.85±1.00 5.85±1.47 4.96
@@
Origin Bb-ar mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 0.3280±0.01 0.2259±0.02 0.3547±0.05 0.3213±0.06 0.2252±0.05 0.2802±0.05 0.2892
U.S. 0.3355±0.02 0.2746±0.05 0.3581±0.03 0.3261±0.05 0.2419±0.03 0.3194±0.02 0.3093
Turkey 0.3141±0.01 0.2434±0.04 0.3154±0.03 0.3111±0.06 0.2210±0.03 0.3164±0.06 0.2869
Japan 0.2876±0.02 0.2411±0.04 0.3150±0.05 0.3030±0.05 0.1834±0.04 0.2553±0.04 0.2642
France 0.3599±0.02 0.2984±0.04 0.3462±0.03 0.3129±0.04 0.2608±0.02 0.2844±0.01 0.3104
U.K. 0.3450±0.02 0.3118±0.08 0.3651±0.04 0.3450±0.06 0.2406±0.07 0.3435±0.05 0.3252
Mexico 0.3098±0.02 0.2891±0.04 0.3335±0.07 0.2687±0.03 0.2163±0.02 0.2665±0.02 0.2807
India 0.2850±0.03 0.2474±0.07 0.3051±0.03 0.3355±0.09 0.2217±0.05 0.2455±0.03 0.2734
Germany 0.3300±0.02 0.2859±0.07 0.3530±0.04 0.2772±0.06 0.2046±0.04 0.3172±0.05 0.2946
China 0.2938±0.02 0.2836±0.06 0.3255±0.07 0.2889±0.03 0.1907±0.04 0.2774±0.04 0.2767
Iran 0.3453±0.04 0.2644±0.06 0.3634±0.03 0.3488±0.04 0.2405±0.06 0.3257±0.02 0.3147
Greece 0.3241±0.02 0.2559±0.04 0.3502±0.05 0.3358±0.06 0.2325±0.03 0.3052±0.04 0.3006
Spain 0.3659±0.02 0.3213±0.03 0.4071±0.06 0.3626±0.06 0.2869±0.05 0.3576±0.05 0.3502
Russia 0.3323±0.02 0.2772±0.02 0.3347±0.02 0.3451±0.02 0.2964±0.03 0.3145±0.00 0.3167
ALL 0.3243±0.01 0.2625±0.03 0.3421±0.04 0.3176±0.05 0.2282±0.04 0.2957±0.03 0.2951